i've tried this questin but i don't get how to do it.
Michelle threw a ball vertically upwards, with its height h, in metres, after a time of t seconds, being given the formula:
h=8t-t^2
Q. find after what time the ball is first at a height of 12m

Found 2 solutions by stanbon, jenrobrody:
Answer by stanbon(48546)

h=8t-t^2
Q. find after what time the ball is first at a height of 12m
Let the height be 12m and solve for "t" as follows:
12=8t-t^2
t^2-8t+12=0
(t-6)(t-2)=0
t=2 seconds is the 1st time the ball is at 12m
Cheers,
Stan H.

Answer by jenrobrody(19)

Let h=12 and solve for t:
12=8t-t^2 set equal to zero(add -8t and t^2 to both sides)
t^2-8t+12=0 can be factored:
(t-6)(t-2)=0 zero product rule:
t-6=0 and t-2=0
t=6 and t=2
The ball is at a height of 12 at times 2 seconds and 6 seconds.
The first time it reaches a height of 12 is 2 seconds.
check:
h=8(2)-2^2
h=16-4=12